# The Signals of the Doomsday

### v3

@@ 7 Modified Higgs particles production due to the friction
We discretize proper time $\tau$ and integrate the coupled system $\{y(\tau),t(\tau),R(\tau),N_{\rm tot}(\tau)\}$ defined by (7.3)–(7.8) using a classical fourth-order Runge–Kutta (RK4) method with adaptive steps. The state is initialized at nucleation by $y(0)=0$ $(v=0)$, $t(0)=0$, $R(0)=R_{0}$, and $N_{\rm tot}(0)=0$. At each substep we evaluate the raw proper acceleration
$\alpha_{\rm raw}(\tau)=A-\frac{2}{R(\tau)}-B\,\sinh y(\tau).$
$\frac{dy}{d\tau}=\alpha_{\rm raw},\qquad\frac{dt}{d\tau}=\cosh y,\qquad\frac{dR}{d\tau}=\sinh y,$
and accumulate the produced quanta using
$\frac{dN_{\rm tot}}{d\tau}=N_{k=0}(\tau)\,4\pi\,R(\tau)^{2}\,\sinh y(\tau),\qquad N_{k=0}(\tau)=\left[\frac{(\omega_{+}+\omega_{-})^{2}}{(\omega_{+}-\omega_{-})^{2}}e^{\frac{4\omega_{+}}{\alpha_{raw}}}-1,\right]^{-1}.$
$N_{k=0}(\tau)\;\simeq\;\left[\frac{(\omega_{+}+\omega_{-})^{2}}{(\omega_{+}-\omega_{-})^{2}}\right]^{-1}e^{-4\mu/\alpha_{\rm raw}}\quad\text{when}\quad\frac{4\mu}{\alpha_{\rm raw}}\gtrsim 200,$
$d\tau\;\sim\;\frac{0.05}{\max\!\big{(}|\alpha_{\rm raw}|,10^{-6}\big{)}},$
Stopping criterion and diagnostics. The integration stops when the coordinate time first satisfies $t(\tau)=\tau_{\rm term}$ for the scenario. For accuracy control, we repeat each run with the initial $d\tau$ halved and demand relative changes $\lesssim 10^{-4}$ in $\{\tau_{\rm term},R_{\rm fin},N_{\rm tot}^{\rm(int)}\}$. As additional consistency checks we verify: (a) the differential inequality $d\alpha/d\tau\leq-B\,\alpha$ (hence $\alpha(\tau)\leq\alpha(0)e^{-B\tau}$) is respected pointwise; (b) monotonicity of $R(\tau)$; and (c) that $N_{k=0}(\tau)$ vanishes whenever $\alpha_{\rm raw}\leq 0$. To illustrate, we consider three ultra-relativistic terminal deficits $\delta\equiv 1-v_{\rm term}$ and use the given friction pairs $(\eta,\tau_{\rm term})$ for each scenario; this fixes $B=\eta/\sigma$ and thus the proper-time dynamics via (7.4)–(7.8). At $t=\tau_{\rm term}$ we report: the final proper time $\tau_{\rm fin}$, radius $R_{\rm fin}$, and the integrated yield $N_{\rm tot}^{\rm(int)}$.
$\gamma_{\rm term}\,v_{\rm term}\;=\;\frac{\Delta V}{\eta},$ (7.9)
$\eta(\delta)\;=\;\frac{\Delta V}{\gamma_{\rm term}\,v_{\rm term}}.$ (7.10)
Scenario $\eta\,[M_{p}^{4}]$ $\tau_{\rm term}\,[M_{p}^{-1}]$ $R_{\rm fin}\,[M_{p}^{-1}]$ $N_{\rm tot}^{\rm(int)}$
$\delta=10^{-8}$ $8.300019459820\mathrm{e}{-16}$ $7.069863\mathrm{e}{+07}$ $7.208700\mathrm{e}{+07}$ $1.052063\mathrm{e}{+08}$
$\delta=10^{-9}$ $624696593960\mathrm{e}{-16} 2.2$$235687\mathrm{e}{+08}$ $2.320814\mathrm{e}{+08}$ $3.530221\mathrm{e}{+09}$
$\delta=10^{-10}$ $8.300019398190\mathrm{e}{-17}$ $7.069863\mathrm{e}{+08}$ $7.147002\mathrm{e}{+08}$ $1.136138\mathrm{e}{+11}$
Table 3: For each ultra-relativistic deficit $\delta=1-v_{\rm term}$ we evolve the spherical wall with proper acceleration (7.4) and kinematics to the time $\tau_{\rm term}=\sigma/\eta$ where the terminal velocity is reached. The proper acceleration $\alpha(\tau)$ directly feeds the instantaneous occupancy (7.6); the total yield $N_{\rm tot}^{\rm(int)}$ follows from (7.8). We note that we took into account only the Higgs particles produced at rest, i.e. the $k=0$ mode. The number of particles produced in all momentum modes are certainly higher than shown here.
The results show that as the radius grows friction balances the drive, with the system approaching terminal velocity on the timescale $\tau_{\rm term}$; and the integrated yield is dominated by the combination of the exponentially sensitive $N_{k=0}(\alpha)$ and the rapidly growing area factor $4\pi R^{2}$, which becomes substantial by the time $\tau\sim\tau_{\rm term}$ in the most ultra-relativistic cases.
These Higgs bosons will decay almost exclusively into Standard Model final states, with the dominant channels producing high‐energy photons and neutrinos. In Figures 6 and 7 below, we present the resulting photon and neutrino spectra, respectively, showing the differential number of particles per event as a function of their energy in Planck‐mass units. These distributions encapsulate both the initial zero‐momentum mode occupation and the volume amplification that occurs during the ultra‐relativistic expansion of each bubble wall. In more intuitive terms, even though only a handful of quanta start life in the bubble wall’s zero mode, the rapid expansion amplifies them into a signal of Higgs decays—potentially producing observable flashes of photons and bursts of neutrinos in a cosmological first‐order transition.
Figure 6: The spectrum of photons generated by true‐vacuum bubbles via Higgs decay. The three curves correspond to Higgs particles produced with different terminal wall velocities, characterized by $\delta=10^{-8}$, $10^{-9}$, and $10^{-10}$, respectively, as summarized in Table 3.
Figure 7: The spectrum of neutrinos generated by true‐vacuum bubbles via Higgs decay. The three curves correspond to Higgs particles produced with different terminal wall velocities, characterized by $\delta=10^{-8}$, $10^{-9}$, and $10^{-10}$, respectively, as summarized in Table 3.
8 Photon or neutrino signal lead time

### v4

@@ 7.1 Thermal particle production
The dynamical system is conveniently summarised as the coupled set
$\left\{\frac{dR}{d\tau},\;\frac{dy}{d\tau},\;\frac{dy_{0}}{d\tau},\;\frac{d\rho_{\rm th}}{d\tau},\;\frac{dN_{\rm th}}{d\tau}\right\},$ (7.27)
which is integrated simultaneously from nucleation up to a finite proper time $\tau_{\rm final}$. Here $(R(\tau),y(\tau))$ describe the physical wall subject to friction, while $y_{0}(\tau)$ is a frictionless reference rapidity evolved at the same instantaneous radius $R(\tau)$ and governed by Eq. (7.9). Evolving both systems in parallel ensures that the energy deficit is always evaluated against the correct frictionless configuration at the same radius.
Astrophysical microphysics is parametrised by the terminal–velocity deficit
$\delta\equiv 1-v_{\rm term},$ (7.28)
which fixes how close the wall comes to the speed of light before friction balances the vacuum pressure. For a given $\delta$ one has
$v_{\rm term}=1-\delta,\qquad\gamma_{\rm term}=\frac{1}{\sqrt{1-v_{\rm term}^{2}}}.$ (7.29)
At terminal velocity the net force on the wall must vanish, so the outward vacuum pressure $\Delta V$ is exactly balanced by the drag (in the planar limit) :
$\Delta V=\eta(0)\,\gamma_{\rm term}v_{\rm term},$ (7.30)
which defines a reference drag coefficient
$\eta(0)=\frac{\Delta V}{\gamma_{\rm term}v_{\rm term}}.$ (7.31)
Smaller $\delta$ corresponds to a more ultra–relativistic wall and hence to a smaller $\eta(0)$; because $\gamma_{\rm term}$ grows extremely rapidly as $v_{\rm term}\to 1$, even modest changes in $\delta$ can change $\eta(0)$ by many orders of magnitude. This motivates our use of $\delta$ as the primary input parameter.
In reality the friction is not constant. As the wall deposits energy into the surrounding medium, a shocked layer builds up and pushes back on the wall with a force that grows with the local temperature. To capture this behaviour we promote the drag coefficient to a temperature–dependent quantity
$\eta(\tau)=g_{\rm eff}^{2}\,T_{\rm eff}^{4}(\tau),$ (7.32)
where $g_{\rm eff}$ is an effective coupling which encodes how efficiently the hot plasma transfers momentum to the wall and $T_{\rm eff}(\tau)$ is an effective temperature combining a small ambient seed temperature with the dynamically generated shock temperature in the thermal shell:
$T_{\rm eff}^{4}(\tau)=T_{\rm amb}^{4}+T_{\rm shock}^{4}(\tau),\qquad T_{\rm shock}^{4}(\tau)=\frac{30}{\pi^{2}g_{*}}\,\rho_{\rm th}(\tau).$ (7.33)
This form is suggested by simple kinetic theory: the drag is proportional to the momentum flux carried by thermal particles, which scales as $n\,p\sim T^{3}\times T\sim T^{4}$, and to the probability $\sim g_{\rm eff}^{2}$ that these particles scatter and transfer their momentum to the wall.
The frictional pressure entering the wall equation of motion is $P_{\rm fric}(\tau)\equiv\eta(\tau)\,\gamma(\tau)v(\tau);$ with the temperature–dependent ansatz $\eta(\tau)=g_{\rm eff}^{2}\,T_{\rm eff}^{4}(\tau)$ this becomes
$P_{\rm fric}(\tau)=g_{\rm eff}^{2}\,\gamma(\tau)v(\tau)\,T_{\rm eff}^{4}(\tau),$ (7.34)
consistent with ultrarelativistic bubble–wall friction calculations in which the thermal backreaction is controlled by the local plasma temperature together with the wall boost (up to coupling and phase–space factors) [81, 65, 13]. Parametrically, this is consistent with detailed calculations of electroweak bubble friction, where the thermal pressure (and thus the effective drag) on an ultra–relativistic wall scales as $P_{\rm th}\propto\gamma^{n}\,T^{4}$ (up to coupling/phase-space factors), with $n$ depending on the dominant kinematics and emission spectrum, see e.g. the all–orders analysis of Ref.[81, 65, 13].
Relativistic scaling of the drag.
Following the ultrarelativistic discussion of Long & Turner[81], the thermal pressure on the wall may be schematically written in the wall frame as $P_{\rm th}\sim F_{a}\langle\Delta p_{z}\rangle$, where the incident flux scales as $F_{a}\sim\gamma\,T^{3}$. For the classic $1\!\to\!1$ kinematics (mass change across the wall with no emission), one has $\Delta p_{z}\propto 1/E_{a}\sim 1/(\gamma\,T)$, so $P_{\rm th}\propto\gamma^{0}$. Including soft/collinear $1\!\to\!2$ emission changes the momentum transfer: for massive emission channels the transfer can be set by the emitted mass scale, yielding $P_{\rm th}\propto\gamma^{1}$. By contrast, if the emission spectrum is approximately log-flat up to a UV cutoff of order the incident energy $E_{a}\sim\gamma\,T$, then $\langle\Delta p_{z}\rangle\sim E_{a}$ and the pressure can scale as $P_{\rm th}\propto\gamma^{2}$.
Motivated by this structure, we adopt the phenomenological drag which explicitly realizes the $\gamma^{1}$ scaling associated with soft/collinear emission in massive channels, while packaging microphysical coupling/phase-space factors into $g_{\rm eff}^{2}$ and the local energy density into $T_{\rm eff}^{4}$. Contributions that can enhance the scaling toward $\gamma^{2}$—e.g. harder (UV-dominated) emission and approximately log-flat spectra extending up to a UV cutoff—are not included in our baseline model. Our quoted thermal yields should therefore be interpreted as conservative in the ultrarelativistic regime.
To connect the dynamic ansatz (7.32) to the terminal condition (7.31), we fix, for each $\delta$, the ambient temperature $T_{\rm amb}$ by demanding that at nucleation ($\rho_{\rm th}=0$, $T_{\rm shock}=0$) the instantaneous drag reproduces $\eta(0)$:
$\eta(0)=g_{\rm eff}^{2}\,T_{\rm eff}^{4}(0)=g_{\rm eff}^{2}\,T_{\rm amb}^{4},$ (7.35)
which yields
$T_{\rm amb}=\left[\frac{\eta(0)}{g_{\rm eff}^{2}}\right]^{1/4}.$ (7.36)
In our numerical Higgs benchmarks we adopt
$g_{\rm eff}=10^{-3},$ (7.37)
[…]
The thermal energy density and multiplicity are updated using the energy–deficit law [^2] Note that while the friction pressure scales as $P_{\rm fric}=\eta\,\gamma v,$ the energy deposition rate entering $d\rho_{\rm th}/d\tau$ scales as work done per unit time, hence carries an additional factor $\sim\gamma v,$ giving a contribution $\propto\eta\,\gamma^{2}v^{2}.$
$\frac{d\rho_{\rm th}}{d\tau}=\gamma\,\mu\,\big(v_{0}-v\big)\left(\Delta V-\frac{2\sigma}{R}\right)+\eta(\tau)\,\mu\gamma^{2}v^{2},$ (7.38)
combined with the usual equilibrium relations
$T(\tau)=\left(\frac{30}{\pi^{2}g_{*}}\,\rho_{\rm th}(\tau)\right)^{1/4},\qquad n_{\rm th}(\tau)=\frac{\zeta(3)}{\pi^{2}}\,g_{*}\,T^{3}(\tau),$ (7.39)
and the comoving production rate
$\frac{dN_{\rm th}}{d\tau}=4\pi R^{2}(\tau)\,\sinh y(\tau)\;n_{\rm th}(\tau).$ (7.40)
Here $\mu$ is the Higgs–like mass scale in the true vacuum and $g_{*}=106.75$.
Integration time. It is convenient to introduce the constant–friction reference timescale
$\tau_{\rm term}=\frac{\sigma}{\eta(0)},$ (7.41)
which is the natural proper time controlling the approach to terminal velocity in the auxiliary system with fixed drag $\eta(0)$. In that simplified model the solution of the linearised equation of motion from Eqn (7.6) in the late time scenario for a large bubble (neglecting the curvature term) shows an approximately exponential approach to the asymptotic value,
$\gamma(\tau)v(\tau)\simeq\gamma_{\rm term}v_{\rm term}\Bigl(1-e^{-\tau/\tau_{\rm term}}\Bigr),$ (7.42)
so that by time $\tau$ a fraction $1-e^{-\tau/\tau_{\rm term}}$ of the asymptotic impulse has been accumulated.
[…]
$\tau_{\rm final}=k\,\tau_{\rm term},$ (7.43)
with
$k=5,\qquad\tau_{\rm final}=5\,\tau_{\rm term}.$ (7.44)
[…]
Finally, we do not extend the integration arbitrarily far beyond $\tau_{\rm final}$, even though in principle an expanding bubble could continue to radiate and produce particles as long as it cruises near its terminal speed. Our estimate is explicitly local and comoving: the shell is moving relativistically, the radiation it emits is boosted and streams out until it leaves the local horizon, and this outgoing flux induces a radiation–reaction (self–force) on the wall. At the same time, ambient particles scatter off the shell, absorb energy, and are kicked away. These dissipative channels provide additional sources of friction, effectively encoded in the last term of Eq. (7.38), and are not contained in the surface tension alone. In practice we find that evolving up to $\tau_{\rm final}=5\,\tau_{\rm term}$ already captures the dominant part of the thermal production: extending the evolution further changes $N_{\rm th}$ only mildly within our present idealised setup, while truly reliable late–time behaviour would require including additional physics — strong gravity, hydrodynamic backreaction of the shocked plasma, and detailed microphysics of radiation reaction and scattering. Our finite $\tau_{\rm final}$ should therefore be viewed as a conservative upper bound on the thermal multiplicity within the regime where our approximations remain under quantitative control.
Numerical results
For the Higgs false–vacuum benchmark we take
$\Delta V=5.869\times 10^{-12}M_{p}^{4},\qquad R_{0}=2\times 10^{4}M_{p}^{-1},\qquad\sigma=5.8680\times 10^{-8}M_{p}^{3},$
with $g_{*}=106.75$ and $\zeta(3)=1.202056$. The true–vacuum Higgs–like mass scale is
$\mu=7.16\times 10^{-4}M_{p},$ (7.45)
which sets the shocked–layer thickness $\ell\sim 1/\mu$ in the heating term. We consider three ultra–relativistic terminal–velocity deficits,
$\delta=10^{-8},\qquad 10^{-9},\qquad 10^{-10},$
compute $\eta(0)$ from Eq. (7.31), fix $T_{\rm amb}(\delta)$ as above, and evolve the coupled system with the dynamic drag coefficient (7.32) up to
$\tau_{\rm final}=5\,\tau_{\rm term},\qquad\tau_{\rm term}\equiv\frac{\sigma}{\eta(0)}.$ (7.46)
For each run we integrate the proper–time equations $dR/d\tau=\sinh y$ and $dy/d\tau=A-2/R-(\eta/\sigma)\sinh y$ with $A\equiv\Delta V/\sigma$, and compute the thermal number from $dN_{\rm th}/d\tau=(dV/d\tau)\,n(T)$ with $dV/d\tau=4\pi R^{2}\sinh y$. The resulting thermal yields are summarised in Table 2.
Scenario ($\delta$) $\eta(0)\,[M_{p}^{4}]$ $\tau_{\rm final}\,[M_{p}^{-1}]$ $N_{\rm th}$
$10^{-8}$ $8.30\times 10^{-16}$ $3.53\times 10^{8}$ $2.15\times 10^{26}$
$10^{-9}$ $2.62\times 10^{-16}$ $1.12\times 10^{9}$ $1.87\times 10^{27}$
$10^{-10}$ $8.30\times 10^{-17}$ $3.53\times 10^{9}$ $1.61\times 10^{28}$
Table 2: Thermal Higgs–channel particle yields obtained from the energy–deficit formulation with a temperature–dependent drag coefficient $\eta(\tau)=g_{\rm eff}^{2}T_{\rm eff}^{4}(\tau)$ and $g_{\rm eff}=10^{-3}$. For each $\delta$, the reference value $\eta(0)$ is fixed by the terminal condition $\Delta V=\eta(0)\gamma_{\rm term}v_{\rm term}$, and the evolution is integrated up to $\tau_{\rm final}=5\,\tau_{\rm term}$.
For fixed $\Delta V$, the terminal matching implies $\eta(0)\propto(\gamma_{\rm term}v_{\rm term})^{-1}$, so $\eta(0)$ decreases as $\delta$ is reduced. In the fully dynamical system this does not imply a smaller thermal yield: weaker friction allows the wall to expand over a larger four–volume before heating shuts off, and the integrated swept volume $\int d\tau\,4\pi R^{2}\sinh y$ grows strongly with decreasing $\delta$. In our benchmarks the thermally produced multiplicities are
$N_{\rm th}\sim 10^{26}\text{--}10^{28},$ (7.47)
which are enormous on particle–physics scales.
If we interpret these quanta as true–vacuum Higgs excitations of mass $m_{H}\simeq\mu=7.16\times 10^{-4}M_{p}$, the total decay energy stored in the hot shell is
$E_{\rm tot}\sim N_{\rm th}\,m_{H}.$ (7.48)
If a fraction $f_{\nu}$ of this energy is ultimately channelled into $\mathcal{O}(1$$100)\,\mathrm{GeV}$ neutrinos (and similarly into photons), the corresponding neutrino multiplicity is roughly
$N_{\nu}\sim f_{\nu}\,N_{\rm th}\,\frac{m_{H}}{\langle E_{\nu}\rangle}\sim 10^{37}\text{--}10^{39}\,\left(\frac{f_{\nu}}{10^{-2}}\right)\left(\frac{10~\mathrm{GeV}}{\langle E_{\nu}\rangle}\right),$ (7.49)
for the range of $N_{\rm th}$ in Table 2. Although this is many orders of magnitude below the neutrino yield of a core–collapse supernova at the source, the absolute number of quanta is still macroscopically large. For a bubble nucleated within $\mathcal{O}(10)\,\mathrm{pc}$ of the Earth such multiplicities correspond to $\mathcal{O}(10^{3}$$10^{7})$ neutrinos crossing a km2–scale detector, so a nearby late–time phase–transition event would appear as an exceptionally bright, short–duration burst in high–energy neutrinos and photons. At cosmological distances, of order $10^{6}$$10^{8}$ light years, the same total energy released in very high–energy photons would manifest as an extremely luminous, pointlike transient — a “shiny” high–energy source on the sky.
7.2 Thermal spectrum of the produced particles
Once the total thermal energy and particle number have been determined from the dynamical evolution, the phase–space distribution of the produced particles is fixed by equilibrium thermodynamics. Since the thermal bath is generated in a thin shocked layer and rapidly equilibrates, the momentum distribution of any light bosonic species is well described by the massive Bose–Einstein spectrum
$f(k)\equiv\frac{k^{2}}{\exp\!\left(\sqrt{k^{2}+m^{2}}/T\right)-1},$ (7.50)
where $m$ is the mass of the scalar produced (for the Higgs-like field used in our benchmark, $\mu=7.16\times 10^{-4}M_{p}$) and $T$ is the thermalisation temperature extracted from the shocked plasma.
Normalisation to the physical yield. The shape of the spectrum is fixed by (7.50), but the overall normalisation must reproduce the total thermal particle number $N_{\rm th}$ obtained from the integrated heating history in Eq. (7.25). We therefore define the physically normalised spectrum
$\frac{{\rm d}N}{{\rm d}k}=N_{\rm th}\;\frac{f(k)}{\displaystyle\int_{0}^{\infty}f(k)\,{\rm d}k},$ (7.51)
which guarantees
$\int_{0}^{\infty}\frac{{\rm d}N}{{\rm d}k}\,{\rm d}k=N_{\rm th}.$ (7.52)
This ensures that the curves shown in Fig. 4 represent the correct physical particle numbers generated by frictional dissipation for each value of the terminal-velocity deficit $\delta$.
Peak structure and scaling. For bosonic thermal distributions with $m\ll T$, the maximum of (7.50) occurs very close to
$k_{\rm peak}\simeq 2.8\,T,$ (7.53)
a standard result that remains accurate for the Higgs-like mass used here. Because smaller $\delta$ leads to much larger thermal energy densities and thus larger temperatures $T$, the peak of the spectrum shifts systematically to higher momenta as $\delta$ decreases. This scaling, together with the large hierarchy in total particle numbers, explains the structure and vertical separation of the three curves in the final spectrum.
Resulting thermal spectrum. Figure 4 displays the physically correct, fully normalised spectrum ${\rm d}N/{\rm d}k$ for the three choices of $\delta$ considered in this work. Each curve encodes: (i) the massive Bose–Einstein shape determined by the shock temperature, (ii) the shift of the peak at $k_{\rm peak}\simeq 2.8T$, and (iii) the correct total particle yield fixed by $N_{\rm th}$. The steep decline at $k\gg T$ reflects Boltzmann suppression in the massive tail of the distribution. Although the plot explicitly shows the spectrum for the Higgs scalar, the same thermal bath also populates all lighter Standard Model degrees of freedom with comparable characteristic momenta $k\sim T$.
Figure 4: Physically normalised thermal spectrum ${\rm d}N/{\rm d}k$ for the Higgs-like scalar with mass $\mu=7.16\times 10^{-4}M_{p}$ produced via frictional dissipation. Each curve corresponds to a different terminal-velocity deficit $\delta$, which determines both the shock temperature and the total thermal particle number $N_{\rm th}$. The peak of each distribution occurs at $k_{\rm peak}\simeq 2.8T$, as expected for a massive Bose–Einstein spectrum. The vertical normalisation of the curves reflects the physical particle yields shown in Table 2.
These results demonstrate that thermalization of the dissipated energy dominates overwhelmingly over vacuum-mismatch particle emission when the terminal-velocity deficit is small. A tiny friction coefficient allows the wall to reach extreme Lorentz factors and sustain a long relaxation period, during which the deficit between the physical and frictionless trajectories is gradually converted into heat. The energy-deficit formulation ensures that this conversion is computed correctly and because the energy-deficit approach tracks the exact portion of wall energy that must be dissipated irrespective of microscopic uncertainties, our analysis yields a robust and model-independent estimate of thermal particle production.
The enormous population of Higgs particles produced through this thermal channel is not stable: each Higgs boson rapidly decays into Standard Model final states, with dominant branching fractions into photons, neutrinos, and charged leptons. Once emitted, the photons free–stream across interstellar and intergalactic distances essentially unimpeded, while the accompanying neutrinos traverse the cosmos with negligible attenuation. Consequently, the Higgs decays generate a distinctive observational signature at Earth in the form of transient gamma-ray bursts and correlated high-energy neutrino bursts. Existing and upcoming gamma-ray detectors and future radio-Cherenkov neutrino arrays—operate in the energy and time windows relevant for short-duration, broadband photon and neutrino bursts of this type. However, because these messengers dilute as $1/d^{2}$ and must compete with astrophysical backgrounds, an observable signal is expected only for bubbles nucleated at relatively short distances on astrophysical scales. By contrast, the violent acceleration and shock heating of the wall also source gravitational radiation whose strain decays more slowly and can remain detectable over cosmological distances. For suitable choices of parameters, the resulting high-frequency burst could fall within the sensitivity reach of space-based interferometers (e.g. LISA) and future facilities such as Cosmic Explorer and the Einstein Telescope. Thus, the thermal production mechanism does not merely represent an internal energy-loss channel of the bubble wall: it implies a concrete, though likely rare, multi-messenger target, in which photons and neutrinos probe sufficiently nearby bubbles while the associated gravitational-wave burst can, in principle, be seen from much farther away, with the overall signal strength directly tied to the ultra-relativistic dynamics of the transition.
7.3 Sub-Planckian temperature regime.
A related conceptual issue is the very high shock temperatures that the model could in principle generate. Even though $\Delta V$ is fixed and local energy conservation is enforced, the total vacuum energy available to the wall is not bounded: as the bubble expands it converts an ever larger volume of false vacuum, so the released energy grows as $\sim\Delta V\,R^{3}$. If all of this work were deposited into an infinitesimally thin shell of fixed thickness, the local thermal energy density would grow roughly as $\rho_{\rm th}\propto\Delta V\,R$ and could formally approach or exceed the Planck scale.
In practice, once $\rho_{\rm th}$ approaches the fundamental high scale of the problem (and in particular $M_{p}^{4}$) our effective field-theory and flat-space descriptions cease to be reliable: strong-gravity effects, gravitational backreaction, and possibly new microscopic degrees of freedom are expected to become important, and one may even encounter scenarios of shell-like gravitational collapse or black–membrane–like configurations[70, 71, 67, 14, 106, 93, 47, 55, 95, 111, 64, 26, 66, 59, 61, 31, 54, 82, 35]. A faithful treatment of such a regime would require including dynamical gravity and going beyond the simple hydrodynamic + thin-wall approximation employed here. [^3] One could refer to these articles for a detailed hydrodynamic treatment relevant to our study. Ref [4, 2, 74, 75, 15, 76, 78, 109, 108, 40, 39, 8]
In the present work we deliberately stay away from this strongly coupled, Planckian regime. With the choice $\ell\simeq 1/\mu$ (corresponding to a comoving layer of thickness $\ell\sim 1/\mu$) and with our benchmark values of $\Delta V$, $\sigma$ and $\delta$, the numerical evolution up to $\tau_{\rm final}=5\,\tau_{\rm term}$ always yields sub-Planckian energy densities and temperatures, $\rho_{\rm th}\ll M_{p}^{4}$ and $T(\tau_{\rm final})\ll M_{p}$. We therefore do not need to impose an explicit hard cutoff in $T$ in the runs shown in Fig. 4: all quoted thermal yields are obtained within a regime where the EFT and flat-space approximations remain under quantitative control.
If one were to explore more extreme regions of parameter space in which the formal evolution drives $T(\tau)$ close to $M_{p}$, the correct description would have to be modified at that point, either by introducing an explicit sub-Planckian cutoff or, more ambitiously, by coupling the wall–plasma system to full gravitational dynamics and possible new high-scale physics. Such Planckian-shell scenarios are intriguing in their own right, but lie beyond the scope of the present analysis and are left for future work.
Taken together, the choices of $\tau_{\rm final}$ and of the comoving layer prescription $\ell\sim 1/\mu$ ensure that our quoted thermal yields constitute a conservative bound computed entirely in a sub-Planckian regime, while avoiding spurious contributions from late-time or super-Planckian dynamics that our simplified framework is not designed to describe.
8 Signal lead time